Solve the system by substitution.
x + 5y = 0
–2x – 4y = 6

Answer by justices98(1)

Solve the system by substitution.
x + 5y = 0
–2x – 4y = 6
|||||||||||||||||||
alright im almost 90% positive ill make the answer right. but anyways
since the bottom equation has negatives you could probly subtract the -4y from
-2x -4y = 6 and add it to the other side making it -2x = 6 + 4y. divide -2 from 6 and 4y making the equation look like this x = -3 - 2y, then SUBSTITUTE x to the top equation. 1(-3 -2y) + 5y = 0, take off the parentheses by multiplying 1 by -3 and -2y making the equation -3 -2y +5y = 0, subtract the y from the other y making
-3 + 3y = 0, then add three to the other side, like this -3 +3 +3y = 0 +3
that gets rid of the 3 on the left side and makes it go on the right side, making
3y = 3, divide and you get y = 1, substitue that for the y's and figure out x then you got a vertex and you're done. gotta let you do some for yourself =]
i hope i did that right but, yea
justices98